OpenStudy (anonymous):

indifinite integral of 1/(2cos^2(x))

zepdrix (zepdrix):

\[\Large\bf\sf \color{royalblue}{\cos x\quad=\quad \frac{1}{\sec x}}\]

OpenStudy (gorv):

1/cosx=secx \[\int\limits_{}^{} \frac{ \sec ^{2} x}{ 2 } *dx\]

OpenStudy (gorv):

\[\frac{ d(tanx) }{ dx }=\sec ^{2}x\]

OpenStudy (gorv):

\[\int\limits_{}^{} \frac{ d(tanx) }{ dx }*dx\]

OpenStudy (gorv):

now solve it